quickstudent:

The hospital provides ambulance service within a 30 mile radius of the hospital. If this service area is represented graphically, with the hospital located at (0, 0), what is the equation that represents the service area? x^2 + y^2 = 900 x^2 + y^2 = 30 (x - 30)^2 + (y - 30)^2 = 0 (x + 30)^2 + (y + 30)^2 = 0

Vocaloid:

Hint: the equation of a circle is: (x-h)^2 + (y-k)^2 = r^2 Where (h,k) is the center of the circle and r is the radius

Allison:

Wouldn't it be answer C, @Vocaloid ?

psirockin3:

I thought it was A.

Allison:

I mean the formula he has shows it right there in C .-.

Hero:

@Allison if you follow the formula @Vocaloid posted, how did you end up concluding with C as your result?

Allison:

The formula was there.

Hero:

(x-h)^2 + (y-k)^2 = r^2 Where (h,k) is the center of the circle and r is the radius .... What does 30 represent in terms of given equation? What does (h,k) represent in the given problem?

Allison:

30 is miles. H is hospital? K is I don't know..?

Hero:



Hero:



Allison:

Okiee?

Hero:

What are the values of h and k in this case?

Allison:

Hospital?

Hero:

If this service area is represented graphically, with the \(\textbf{hospital}\) located at \(\textbf{(0, 0)}\), what is the equation that represents the service area?

Allison:

Math Processing Error is all I see.

Hero:

@Ultrilliam

Hero:

1496429706-59318823129d6eca295e3100-Hospital.JPG

Hero:

There

Allison:

A?

Ultrilliam:

you could have bolded it with the bold command ya know... and it loaded fine for me ._. sometimes that happens ._.

Ultrilliam:

Happened on OS from time to time to

Ultrilliam:

where LaTeX would break because they published a update to it, and they couldn't do a thing but wait.

Ultrilliam:

(for the people who update latex to fix it)

Hero:

I don't think this particular issue is a LaTeX issue.

Hero:

A is correct @Allison. Thanks for taking the time to answer this question. Hope you enjoy your summer.

Allison:

Thanks. You too.